OpenStudy (anonymous):

A particle travels in the xy-plane so that the position is given by x(t) = sin(t) and y(t) = e-3t - 1. What is the total distance traveled by the particle for 0 ≤ t ≤ 3?

OpenStudy (anonymous):

i think it use integral

OpenStudy (anonymous):

yeah i know the formula is total distance is teh integral of velocity from 0 to 3 but idk what to do next

OpenStudy (anonymous):

\[\Delta x= \int\limits\limits_{0}^{3}\sin (t)\] \[=\left[ -\cos(t) \right] \] =-cos(3)-(-cos(0))

OpenStudy (anonymous):

\[\Delta y=\int\limits_{0}^{3}e-3t-1\] \[=et-3/2 t ^{2}-t\] ={e*3-(3/2)*3^2-3}-{0}

OpenStudy (anonymous):

e=2.718

OpenStudy (anonymous):

\[ \int_0^3 \sqrt{ x'(t)^2 + y'[t]^2} dt \]

OpenStudy (anonymous):

Is \[ y(t) =e ^{-3t} -1 \] or \[ y(t) =e ^{-3t -1} \]

OpenStudy (anonymous):

In either case you end up with a messy integral that needs some numerical approximation.